# Statistical closures from the Martin, Siggia and Rose approach to turbulence

### v1

@@ Abstract
The goal of this paper is to study the statistical closures suggested by the Martin-Siggia and Rose approach to statistical turbulence. We find that the formalism leads to a Bethe-Salpeter equation for the three point correlation of the velocity field. In the leading order approximation this equation becomes an explicit expression, which reproduces the closure found in W D McComb and S R Yoffe, A formal derivation of the local energy transfer (LET) theory of homogeneous turbulence, J. Phys. A: Math. Theor. 50, 375501 (2017). This suggests ways to improve upon this closure by resumming the perturbative expansion and/or applying renormalization group techniques.
@@ I Introduction
The goal of this paper is to discuss the solution of the closure problem in the field theory approach to statistical turbulence [1, 2, 3, 4], more specifcically within the Martin - Siggia and Rose (MSR) framework [5, 6, 7, 8, 9, 10]. We shall derive a Bethe-Salpeter equation [11, 12, 13, 14] for the three point correlation of the turbulent velocity. In the simplest (one-loop) approximation, this equation reduces to an explicit formula, which reproduces the closure advanced by McComb and Yoffe in ref. [15].
@@
A key point in the analysis of ref. [15] is to formulate the problem in such a way that the full correlation function eq. (3) appears as an unknown, rather than performing a perturbative expansion for it. To replicate this within the MSR approach we must go beyond the 1-particle irreducible (1PI) effective action (EA) (see [23, 24, 25, 26, 27]). In this work we shall restrict ourselves to the 2-particle irreducible (2PI) EA [23, 28, 29, 30, 31]; for a discussion of 2PI techniques in a MSR framework see [32, 33]. We leave the consideration of higher irreducible EA’s for future work [34, 35, 36, 37, 38, 39, 40].
We shall show that the 2PIEA leads to a Bethe-Salpeter equation for the three point correlation, which in the simplest approximation, namely one-loop order, reduces to the closure derived in ref [15]. We believe this result validates the functional MSR approach to hidrodynamic turbulence, and opens up interesting new perspectives to be further investigated in future work.
The paper is organized as follows. In the next section II we present the 2PI MSR effective action; for a full discussion see. Then in section III we use it to derive a Bethe-Salpeter equation for the three point correlations, and in section IV we show that to leading order the Bethe-Salpeter equation reduces to McComb and Yoffe’s closure from ref. We conclude with some brief final remarks in section V.
@@ III Bethe-Salpeter equation for three-point correlations
Observe that the left hand side is just the expectation value . This may be contrasted to Novikov’s formula [21, 22]
so
On the other hand, let be any time such that . Then the solution of the NSE 1 for is uniquely determined by the velocity field at and the sources for , which are independent of the sources at . Therefore, we may write
In the limit the second derivative may be computed from the continuum limit of eq. (8)
no longer a random variable, and so we find that
also.
@@
IV Derivation of the McComb-Yoffe closure
In the simplest approximation where we neglect we get
@@ IV Derivation of the McComb-Yoffe closure
which is the closure found in [15].
V Final Remarks

### v2

@@ Abstract
The goal of this paper is to study the statistical closures suggested by the Martin-Siggia and Rose approach to statistical turbulence. We find that the formalism leads to a Bethe-Salpeter equation for the three point correlation of the velocity field. In the leading order approximation this equation becomes an explicit expression. We discuss under which approximations this closure reduces to that proposed in W D McComb and S R Yoffe, A formal derivation of the local energy transfer (LET) theory of homogeneous turbulence, J. Phys. A: Math. Theor. 50, 375501 (2017). This suggests ways to improve upon this closure by dropping these restrictions, resumming the perturbative expansion and/or applying renormalization group techniques.
@@ I Introduction
The goal of this paper is to discuss the solution of the closure problem in the field theory approach to statistical turbulence [1, 2, 3, 4], more specifically within the Martin - Siggia and Rose (MSR) framework [5, 6, 7, 8, 9, 10]. As shown by Krommes [11, 12, 13], the MSR approach leads to a Bethe-Salpeter equation [14, 15, 16, 17] for the three point correlation of the turbulent velocity. In the simplest (one-loop) approximation, this equation reduces to an explicit formula, which under further approximations reproduces the closure advanced by McComb and Yoffe in ref. [18].
@@
A key point in the analysis of ref. [18] is to formulate the problem in such a way that the full correlation function eq. (3) appears as an unknown, rather than performing a perturbative expansion for it. To replicate this within the MSR approach we must go beyond the 1-particle irreducible (1PI) effective action (EA) (see [26, 27, 28, 29, 30]). In this work we shall restrict ourselves to the 2-particle irreducible (2PI) EA, introduced in [31, 32, 33, 34], see also [26, 35, 36, 37, 38]; for a discussion of 2PI techniques in a MSR framework see [11, 39, 40]. We leave the consideration of higher irreducible EA’s for future work [41, 42, 43, 44, 45, 46, 47].
Concretely, our goal is to compare the closure relations obtained from the Bethe-Salpeter equation derived from MSR [11] to the closure relation advanced in ref. [18]. In [18] the problem is studied in the framework of the Fokker-Planck equation for the probability density function (pdf) of equal-time velocity fluctuations, introduced by Hopf [48]. A perturbation scheme is proposed where at “zeroth order” the pdf is assumed to be Gaussian, and then a “first order” closure relation is derived. Our goal is to find out what is the equivalent approximation scheme within the MSR approach. The relationship between the Hopf and MSR approaches is reviewed in [30].
The paper is organized as follows. In the next section II we present the 2PI MSR effective action; for a full discussion see [11, 26, 39]. In section III we review the derivation of the Bethe-Salpeter equation for the three point correlations [11], and in section IV we derive the closure approximation which follows from a first order truncation of the Bethe-Salpeter equation. In section V we compare this closure to the one proposed by McComb and Yoffe’s in ref [18]. We find the necessary assumptions which are the correlate of the assumption of near Gaussianity of the Hopf pdf. Under these assumptions, both closure relations are equivalent. We conclude with some brief final remarks in section VI.
In the Appendix we show a concrete example where the identity between the derivatives of the velocity with respect of the stirring forces and with respect to the initial velocity field may be stablished by an independent argument from the formal proof presented in section V.
@@ III Bethe-Salpeter equation for three-point correlations
IV Lowest order MSR closure
In the simplest approximation where we neglect eq. (65) yields
@@ IV Lowest order MSR closure
which is the lowest order closure from MSR.
V MSR and LET closures, a comparison
The closure given in eq. (4.55) of ref. [18] is identical to the one given in eq. (93), except that, instead of the function defined in eq. (87), a new function appears. The new function comes from the Fourier transform
Moreover, the new function only needs to be known to zeroeth order in the LET perturbative scheme. To this order equal time velocity fluctuations are assumed to be Gaussian, and then from Novikov’s formula [24, 25] we obtain the so-called “fluctuation-response relation”, eq. (4.68) in ref. [18],
We emphasize that eq. (95) only holds when the pdf of equal time velocity fluctuations is Gaussian, or under the even more restrictive assumption that the evolution is linear. We shall assume the former.
To prove the equivalence of the closure eq.(93) to the LET closure we must show that actually the two functions and are the same for all times , and then provide the supplementary assumptions under which we recover a MSR “fluctuation-response relation”
These supplementary assumptions are the MSR counterpart of the near Gaussianity assumed in LET.
To show that when we go back to eq. (38) above. Observe that the left hand side of eq. (38) is just the expectation value . This may be contrasted to Novikov’s formula [24, 25]
so
And then, taking the Fourier transform from eq. (87), we get
We have shown that is also related to a variational derivative of the velocity field, but in this case the derivative is taken with respect to the stirring forces, rather than the initial velocities. To complete our argument, we must show that the two derivatives are the same when the two time arguments are different. The proof depends upon the NSE being of first order with respect to time, and the stirring noise being aditive (see Appendix).
Recall that we are assuming that for given initial conditions at time , say, and a noise realization at all times between and , the solution to the NSE between and is unique, which requires Random Galilean Invariance to be broken. Of course, this is also assumed in LET, because otherwise the function would be undefined.
Let be an intermediate time, . Then the solution of the NSE 1 from to is uniquely determined by the velocity field at and the sources for times later than , which are independent of the sources at times earlier than . Therefore, we may write
In the limit the derivative with respect to the driving force may be computed from the continuum limit of eq. (8). Observe that for all times because of causality. Taking the derivative of eq. (8) with respect to the stirring force, and integrating from any negative time to , we get
no longer a random variable, and so we find that
as we wanted to show.
Let us mention that is discontinuos at the coincidence limit, since
If we define , then everywhere.
We now turn to the discussion of the fluctuation-response relation eq. (96). Within the MSR approach we have the exact formula, derived from eq. (41) above (assuming for simplicity)
where is the dressed noise self-correlation, which comes from
For this to reduce to eq. (96) two conditions must be met, first
and second
Since is symmetric, eq. (106) implies that for any . Thus the dressed noise must be white in time, same as the bare noise in eq. (1).
With respect to the composition rule eq. (107), the propagator obeys an equation, also derived from eq. (41) above, of the form
where is the self-energy
Since we are assuming strict inequality the right hand side vanishes. Then eq. (107) requires
or else
Once again, this means that for any , and so the evolution of the propagator must be Markovian
in which case
clearly satisfies the composition rule eq. (107).
In summary, if the dressed noise is nearly white in time and the evolution of the propagator is nearly Markovian, then we obtain the fluctuation-response relation eq. (96) as a zeroth order approximation, and the proof of equivalence of the MSR and LET closures is complete. These assumptions are very strong, but not stronger than the assumption of near Gaussianity in the equal-time velocity fluctuations in the LET approach.
VI Final Remarks
@@
Appendix A G is equal to R, an example
We would like to complement the formal arguments to the effect that the two functions and with a concrete example. Consider the one-dimensional noisy Burgers equation [60, 61, 62]
where the noise is Gaussian and white in time. We wish to compare
to
The Cole-Hopf transformation
transforms eq. (114) into (assuming trivial boundary conditions in spatial infinity)
where
We may transform eq. (118) into an integral equation using the Green function
whereby, given
We now have
and
Now
On the other hand
To compute the second term, we observe that
so
We finally compute the first term in eq. (125)
This establishes the identity of and , since they share both the same equation and the same initial data.
@@ Acknowledgements.
- [11] J. A. Krommes, Turbulence, clumps and the Bethe-Salpeter equation (IAEA-SMR-32/10), in Theoretical and Computational Plasma Physics, International Atomic Energy Agency, Vienna, p. 405 (1978).
- [12] J. A. Krommes, Systematic statistical theories of plasma turbulence and interrnittency: current status and future prospects, Physics Reports 283, 5 (1997).
- [13] J. A. Krommes, Fundamental statistical descriptions of plasma turbulence in magnetic fields, Physics Reports 360, 1 (2002).
@@
- [31] J. Luttinger and J. Ward, Ground-state energy of many-fermion systems, Phys. Rev. 118, 1417 (1960).
- [32] C. de Dominicis and P. C. Martin, Stationary entropy principle and renormalization in normal and superfluid systems II, J. Math. Phys. 5, 31 (1964).
- [33] H.D. Dahmen and G. Jona Lasino, Variational Formulation of Quantum Field Theory, Il Nuovo Cimento LIIA, 807 (1967).
- [34] J.M. Cornwall, R. Jackiw and E. Tomboulis, Effective action for composite operators, Phys. Rev. D 10, 2428 (1974).
@@
- [48] E. Hopf, Statistical hydromechanics and functional calculus, J. Rat. Mech. Anal. 1, 87-123 (1952).
@@
- [60] J. Burgers, The nonlinear diffusion equation, Springer (1974).
- [61] L. Bertini, N. Cancrini, G. Jona-Lasinio, The Stochastic Burgers Equation, Commun. Math. Phys. 165, 211-232 (1994).
- [62] H. Fogedby, Aspects of the noisy Burgers equation, in Anomalous Diffusion From Basics to Applications, Lecture Notes in Physics, Volume 519. Springer-Verlag (1999), p. 101.